OpenStudy (anonymous):

What is the solution to the equation? 8x + 4 = –20 A. x = –3 B. x = –2 C. x = 2 D. x = 3

OpenStudy (anonymous):

anyone understand this?

OpenStudy (anonymous):

x = -3

OpenStudy (anonymous):

how can you tell

OpenStudy (haseeb96):

A =-3 8x+4=-20 8x=-20-4 8x=-24 x=-24\8 x=-3

OpenStudy (anonymous):

Step 1 : Pulling out like terms : 1.1 Pull out like factors: 8x + 24 = 8•(x + 3) Equation at the end of step 1 : 8 • (x + 3) = 0 Step 2 : Solve 8•(x+3) = 0 Equations which are never true : 2.1 Solve 8 = 0 This equation has no solution(s). A a non-zero constant never equals zero. Solving a Single Variable Equation : 2.2 Solve : x+3 = 0 Subtract 3 from both sides of the equation: x = -3

OpenStudy (anonymous):

X= -3 if you plug in -3 where the x is you can solve the equation. 8(-3)= -24 -24+4= -20

OpenStudy (anonymous):

get it now.

OpenStudy (anonymous):

ok thanks, I like it better when people dont just give me the answer, but actually explain it

OpenStudy (anonymous):

ok that's fine welcome

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

i'll invite you ok

OpenStudy (anonymous):

ok... no who should i give a medal to

OpenStudy (anonymous):

i invited you

OpenStudy (anonymous):

can you go